The problem is 2/x^-4. I know the way to solve normal negative exponents so I tried that and got: 2/1/x^4, simplified to 2/x^4 because 2/1 is 2. I got 2/x^4 as a final answer. Is this correct?

Found 2 solutions by oscargut, stanbon:
Answer by oscargut(686)

2/(x^-4) = 2/(1/x^4) = 2* (x^4/1)= 2*x^4 = 2x^4

Answer by stanbon(48535)

The problem is 2/x^-4.
--------------------------
= 2/[1/x^4]
Now you need to invert the denominator and multiply:
= 2
= 2x^4
============
Cheers,
Stan H.